2 moles of water at 10 degree celsius changes to ice at 0 degree celsius. entropy change of water is?

Grade:12

1 Answers

Aarti Gupta
10 years ago
Given-
no. of moles= 2
T1= 10 degree C or 283K
T2= 0 degree C or 273K
del S= ?
sol-
we know that,
del S = delfus H/Tf since, water---->ice so delfus H and Tf
now, delfus H =del ng*R*del T = 2*8.314* (283-273) = 166.28
thus,
del S = 166.28/273 = 0.609 J/K/mol